# Diversity-Enhanced Reasoning for Subjective Questions

### v1

@@ 5 Experiment
Zero-shot CoT 62.45 56.02 32.62 65.88 50.30 52.22 51.82 36.14 63.06 83.83 80.48 68.08
Self-Refine 74.08 73.13 43.13 59.88 50.76 66.09 52.19 37.36 54.02 77.61 87.01 80.37
Role-Play 73.61 74.68 41.67 77.75 52.69 67.43 50.83 37.89 55.07 76.20 85.66 72.87
More think 80.76 80.44 36.42 86.90 60.45 78.82 64.44 81.53 64.50 85.85 82.05 81.79
Ours SelfConsis SFT 85.88 81.67 43.13 85.58 67.35 78.94 67.45 82.19 66.88 83.10 80.62 74.87
Ours SelfConsis SFT+DPO 86.41 60.43 44.20 61.17 68.19 64.09 67.28 68.51 67.24 69.83 81.51 67.56
Ours SelfConsis SFT+GRPO 94.30 85.52 47.22 87.46 70.83 82.15 69.50 85.40 69.43 86.85 85.58 82.16
MultiRole-R1 SelfConsis SFT+GRPO(RS) 94.50 86.25 49.10 89.67 70.85 83.31 66.83 87.27 66.94 87.96 87.36 82.46
@@
Zero-shot CoT 80.89 79.92 38.41 87.07 60.84 73.98 62.46 79.44 67.21 83.39 78.87 76.52
Self-Refine 74.20 75.85 43.19 81.11 61.95 78.87 60.96 80.17 63.77 82.61 80.95 81.24
Role-Play 74.40 80.91 44.87 83.02 62.70 77.32 64.24 79.78 67.32 82.27 77.33 75.02
More think 88.20 84.11 44.04 87.19 64.41 80.30 68.06 83.99 70.42 84.73 83.30 84.12
Ours SelfConsis SFT 89.69 82.64 48.17 87.26 70.05 79.77 70.56 81.36 70.86 83.88 86.02 81.53
Ours SelfConsis SFT+DPO 90.52 80.43 48.89 80.97 69.81 76.62 71.22 83.64 71.28 82.71 86.34 81.81
Ours SelfConsis SFT+GRPO 94.47 85.75 48.55 89.36 69.26 83.37 75.63 87.89 73.71 87.96 87.49 85.31
MultiRole-R1 SelfConsis SFT+GRPO(RS) 95.55 89.58 49.06 91.78 71.48 90.55 75.84 96.54 75.12 92.98 89.79 88.45
@@
Zero-shot CoT 85.01 68.06 36.82 79.18 75.05 71.83 73.63 72.20 75.85 83.09 85.58 70.68
Self-Refine 90.42 80.13 49.04 69.40 71.28 78.22 76.48 83.22 76.55 82.31 84.73 80.24
Role-Play 91.18 81.87 49.90 75.73 67.41 70.74 77.16 75.30 75.71 79.05 91.50 76.60
More think 94.57 80.67 41.60 84.04 75.90 76.81 79.36 83.33 79.36 81.77 88.76 80.94
Ours SelfConsis SFT 94.40 75.06 50.98 81.04 76.08 73.65 81.45 71.34 81.50 77.60 91.61 91.62
Ours SelfConsis SFT+DPO 94.98 67.21 51.33 65.03 75.82 71.71 81.88 42.32 79.82 69.41 90.92 68.69
Ours SelfConsis SFT+GRPO 95.98 86.88 51.73 90.33 75.65 84.92 83.50 89.42 81.19 89.64 91.87 86.36
MultiRole-R1 SelfConsis SFT+GRPO(RS) 97.50 90.17 53.98 91.32 76.50 89.08 86.00 92.89 82.00 91.61 93.43 87.24
@@
Zero-shot CoT 91.71 70.10 42.13 60.99 73.40 57.43 72.29 76.68 80.81 57.84 85.41 70.49
Self-Refine 88.93 53.07 45.25 85.00 69.40 48.16 70.64 49.60 69.22 50.99 84.91 82.31
Role-Play 89.77 41.58 47.57 54.89 70.26 50.49 70.67 47.83 72.98 48.17 93.58 81.93
More think 95.18 74.20 43.39 78.98 75.10 68.44 78.26 72.45 81.20 73.07 90.02 75.07
Ours SelfConsis SFT 94.05 74.02 50.32 77.07 75.96 72.78 78.39 68.35 81.00 73.50 91.62 70.23
Ours SelfConsis SFT+DPO 94.21 74.83 50.68 74.14 76.10 69.58 78.09 50.31 80.45 64.87 91.84 65.32
Ours SelfConsis SFT+GRPO 95.91 85.47 51.37 87.74 77.83 80.36 79.82 86.84 81.19 86.40 91.97 85.98
MultiRole-R1 SelfConsis SFT+GRPO(RS) 96.98 88.15 51.72 89.88 77.95 83.84 81.95 89.09 82.10 87.93 94.98 86.93
Table 1: Main results of the baselines (specified in Section 5) and our proposed method. Acc. is the accuracy of the task (in %) and Div. measures the length normalized diversity score of the reasoning chain (in %). We include two ablations of MultiRole-R1, including SFT on self-consistency filtered data only (Ours SelfConsis SFT), and also SFT with vanilla GRPO (Ours SelfConsis SFT + GRPO). “GRPO(RS)” represents GRPO with reward shaping, which is used in MultiRole-R1. O.O.D. denotes the datasets that are for testing only. Detailed decomposition of the diversity score is shown in Appendix E.
@@ 5.1 Datasets
We train our model on three subjective tasks: ambiguous question answering (BBQ by Parrish et al. (2022)), opinion-based QA (GlobalOpinionQA by DURMUS et al. (2024)), and ethical dilemma (ETHICS by Hendrycks et al. (2021)). To evaluate the effectiveness and generalizability of our approach, we test on three additional datasets: cultural natural language inference (CALI by Huang and Yang (2023)), commonsense reasoning (CSQA by Talmor et al. (2019)), and mathematical reasoning (GSM8K by Cobbe et al. (2021)). Specifically, for the testing data, CALI consists of the out-of-domain (O.O.D.) subjective questions, while CSQA and GSM8K consist of O.O.D. objective questions. It is worth noting that among these benchmarks, GLOQA and CALI have role-dependent ground truths, whereas the others rely on a single ground truth for all roles.
@@ 5.2 Baselines
We first incorporate the following in-context learning Brown et al. (2020) settings: (1) Zero-Shot CoT Kojima et al. (2023); Wei et al. (2023), (2) Role Playing Prompting Kong et al. (2024) and (3) Self-Refine Prompting Madaan et al. (2023).
@@
As observed by Muennighoff et al. (2025), extending the reasoning chain length can further enhance the reasoning capabilities of o1-style models. In MultiRole-R1, this is achieved by suppressing the end-of-thinking token and appending a continuation string (e.g., “wait, I need to think from {role}’s perspective") to encourage extended reasoning from a different role perspective. In the more think baseline, we employ a reasoning length three times longer than regular think, as it offers a balance between efficiency and accuracy based on our pilot analysis.
@@
We perform supervised finetuning on the base model on the self-consistency filtered dataset of size 2,700. Our SFT training and evaluation are conducted via Llama-Factory Zheng et al. (2024).
@@
We introduce another SFT+RL pipeline as an ablation to MultiRole-R1. In this setting, DPO is applied to the self-consistent SFT model, using ground-truth-hinted role answers as positive samples and inconsistent role answers as negative samples.
@@ 5.3 Models
Our experiment is performed on a range of open-source LRMs including R1-Distill-Qwen-7B, R1-Distill-Llama-8B and R1-Distill-Qwen-14B models DeepSeek-AI et al. (2025), and Qwen3-8B Qwen Team (2025) with reasoning mode.
Figure 4: (a) We plot the accuracy against the diversity for each setting (e.g. zero-shot, role-play, morethink, etc.) for each dataset. We observe a positive linear correlation between the two metrics.  (b) Qualitative example of the 34 most frequent roles in the training data generated by LRMs. A more detailed visualization is presented in Figure 5.
@@ 5.4 Metrics
Taking into account the subjective nature of role-based reasoning where the ground truth for subjective questions may vary across different roles, we adopt two different perspective merging strategies during evaluation.
(1) Divergent Merging: for tasks such as CALI and GLOQA, each role $i$’s answer $a_{i}$ is compared with the corresponding ground truth $g_{i}$, where the divergent accuracy is given by:
$Acc_{div}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[a_{i}=g_{i}].$
(2) Convergent Merging: datasets like BBQ, ETHICS, CSQA and GSM8K have answers invariant with role-perspectives. We aggregate different role’s answer to obtain a consensus, and then compare it to the ground truth:
$\hat{a}=argmax\sum_{i}\mathbbm{1}(a_{i}=\hat{a}),Acc_{con}=1[\hat{a}=g].$
@@
To quantify the diversity of model-generated reasoning, we design a composite metric that captures linguistic variation across multiple dimensions. Inspired by prior work on lexical and entropy-based diversity in natural language generation Li et al. (2016); Tanaka-Ishii and Aihara (2015), our metric is a weighted sum of eight complementary diversity signals, including lexical, token entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, distinct N-gram and function word diversity. Formal definition of the diversity metrics can be found in Appendix E. Formally, we express the final combined diversity score in the formula below:
$D_{final}=0.15\sum D_{i}+0.1\sum D_{j},$ (4)
where: $D_{i}\in\{D_{lex},D_{ent},D_{pat},D_{bi}\}$ and $D_{j}\in\{D_{len},D_{adj},D_{yule},D_{func}\}$. The choice of the weighting is further illustrated in Section 7.2.
6 Results and Analysis
@@ 6.1 Main Results
Table 1 shows that with multi-role SFT and GRPO with reward shaping, MultiRole-R1 achieves superior performance in both subjective and objective reasoning tasks, surpassing More Think by an average of 7.6% and 3.8% in accuracy and diversity. Notably, by training solely on subjective tasks, we enhance both the subjective and objective reasoning abilities of models.
@@
For the choice of RL algorithm, we found that GRPO yields more accurate and diverse responses than DPO. We attribute this to a fundamental mismatch between DPO’s training format and the nature of our task. Subjective questions do not have single "correct" answers, yet DPO requires a rigid "preferred-rejected" pair for training. These results therefore demonstrate the superiority of GRPO for this context and validate the effectiveness of explicitly optimizing for diversity.
Interestingly, Table 9 to 12 in the appendix show that our approach results in a decrease in answer tokens compared to SFT, suggesting that mere verbosity does not lead to more accurate answers.
6.2 Domain Generalization
As shown in Table 1, besides in-domain subjective questions that are included in the training set, MultiRole-R1 successfully boosts the performance in all the benchmarks. This shows our approach’s generalizability is two-fold: generalizability to other subjective questions, and also generalizability to objective questions. We reveal that by only training on three subjective datasets, the model gains a generalized understanding of role generation and role reasoning, and has a near 10% accuracy gain in subjective tasks like cultural natural language inference in CALI. Moreover, without any training in objective questions, MultiRole-R1 significantly enhances the models’ performance on commonsense reasoning and math reasoning benchmarks. This demonstrates that subjective question training can potentially enhance the performance of objective questions. An explanation is that the nature of subjective questions pushes the model to gain insights from various perspectives, which collectively guides the model to approach the optimal solution.
Model BBQ GLOQA CALI ETHICS CSQA GSM8K
R1-Distill-Qwen-7B 94.2 45.5 89.2 98.6 92.4 58.4
R1-Distill-Llama-8B 89.0 48.0 82.4 83.6 85.6 64.9
R1-Distill-Qwen-14B 82.0 59.0 60.5 77.2 57.4 79.2
Qwen3-8B 76.0 65.4 89.3 73.0 82.5 33.4
@@ 6.2 Domain Generalization
Table 2: The task break-down correlation coefficient (in %) between accuracy and diversity.
6.3 Accuracy-Diversity Correlations
@@
Our analysis reveals a positive correlation between accuracy and diversity (Figure 4a). This relationship is reinforced by a strong per-task correlation between diversity and accuracy (0.9), which significantly exceeds the correlation with response length (Table 2). These results suggest that performance gains in modern reasoning models like o1 or R1 are driven by a scaling law of diversity, rather than superficial verbosity. Crucially, this demonstrates that optimizing for diversity does not compromise correctness but rather enhances accuracy. One possible explanation is that optimizing for diversity can serve as a useful inductive bias, enabling the model to explore a broader solution space and discover more accurate, perspective-aligned answers in subjective tasks.
Model BBQ GLOQA CALI (O.O.D) ETHICS CSQA (O.O.D) GSM8K (O.O.D) Alignment
@@
6.4 Role Coverage
We leverage the LLM itself to generate roles pertinent to the context of the question. In total, there are 968 distinct roles in the training data, enhancing the perspective diversity of the LRM. We leverage the LLM itself to generate roles relevant to the context of each question. These roles demonstrate broad coverage, including different moral philosophies, nationalities, identity groups, and specific individuals pertinent to the questions. Figure 4 (b) shows a qualitative example of the most frequently occurred roles.
7 Discussion
7.1 Filtering Strategy of the Training Data
Our methodology compares two data sampling strategies for Supervised Fine-Tuning (SFT): an unsupervised self-consistency approach and a supervised method that leverages ground-truth agreement. Table 4 reports the test performance on equal-sized datasets generated by each strategy. The results indicate that while self-consistency filtering can yield slightly lower accuracy in some cases, its performance is broadly comparable to, and can even outperform, supervised filtering. This highlights that using noisy, unsupervised data is sufficient and viable for our tasks Wang et al. (2025d).
BBQ GLOQA CALI ETHICS CSQA GSM8K
(R1-Distill-Qwen-7B)
Consis. Filter 85.55 47.13 67.35 67.45 66.88 80.62
GT Filter 88.40 45.55 65.95 68.44 68.45 80.63
(R1-Distill-Llama-8B)
Consis. Filter 89.69 48.17 72.05 70.56 70.86 83.30
GT Filter 87.44 49.29 69.27 72.15 71.06 84.20
(R1-Distill-Qwen-14B)
Consis. Filter 94.40 50.98 76.98 81.45 80.50 91.61
GT Filter 94.88 52.29 76.28 81.57 80.79 91.28
(Qwen3-8B)
Consis. Filter 94.05 50.32 75.96 78.39 81.00 91.62
GT Filter 94.80 51.07 76.15 80.19 82.13 91.85
7.2 Design of the Diversity Weighting
In equation (4), we design our diversity metric as a combined score of lexical diversity, token entropy diversity, etc. Our guiding principle of diversity metric design is to balance the weights of different diversity factors and make it align with human intuition as much as possible. We present Table 3 showing qualitative alignment scores between the metrics and human diversity ratings. We manually label 60 outputs from different settings, from a diversity scale of 1 to 10. Taking the human rating as a reference, we set the weight of the diversity metrics such that the correlation coefficient between the overall diversity and the human rating is highly aligned. Table 3 shows that our diversity metric has a high alignment score with human ratings, showcasing the reasonableness of our weighting.
7.3 Advanced Math Reasoning
To further validate the robustness of our pipeline on unseen objective tasks, we introduced the AIME 2024 benchmark, which is considerably more challenging than GSM8K. Notably, despite being trained exclusively on subjective datasets, our method demonstrates a consistent O.O.D. performance improvement on AIME, similar to the gains observed on GSM8K.
Model r1-dist-qwen7B r1-dist-llama8B r1-dist-qwen14B qwen3-8B

### v3

@@ 4 Experiment and Results
Zero-shot CoT 62.45 56.02 32.62 65.88 51.82 36.14 50.30 52.22 63.06 83.83 80.48 68.08
Self-Refine 74.08 73.13 43.13 59.88 52.19 37.36 50.76 66.09 54.02 77.61 87.01 80.37
Role-Play 73.61 74.68 41.67 77.75 50.83 37.89 52.69 67.43 55.07 76.20 85.66 72.87
More think 80.76 80.44 36.42 86.90 64.44 81.53 60.45 78.82 64.50 85.85 82.05 81.79
SelfConsis SFT 85.88 81.67 43.13 85.58 67.45 82.19 67.35 78.94 66.88 83.10 80.62 74.87
SelfConsis SFT+DPO 86.41 60.43 44.20 61.17 67.28 68.51 68.19 64.09 67.24 69.83 81.51 67.56
SelfConsis SFT+GRPO 94.30 85.52 47.22 87.46 69.50 85.40 70.83 82.15 69.43 86.85 85.58 82.16
MultiRole-R1 SelfConsis SFT+GRPO(RS) 94.50 86.25 49.10 89.67 66.83 87.27 70.85 83.31 66.94 87.96 87.36 82.46
@@
Zero-shot CoT 80.89 79.92 38.41 87.07 62.46 79.44 60.84 73.98 67.21 83.39 78.87 76.52
Self-Refine 74.20 75.85 43.19 81.11 60.96 80.17 61.95 78.87 63.77 82.61 80.95 81.24
Role-Play 74.40 80.91 44.87 83.02 64.24 79.78 62.70 77.32 67.32 82.27 77.33 75.02
More think 88.20 84.11 44.04 87.19 68.06 83.99 64.41 80.30 70.42 84.73 83.30 84.12
SelfConsis SFT 89.69 82.64 48.17 87.26 70.56 81.36 70.05 79.77 70.86 83.88 86.02 81.53
SelfConsis SFT+DPO 90.52 80.43 48.89 80.97 71.22 83.64 69.81 76.62 71.28 82.71 86.34 81.81
SelfConsis SFT+GRPO 94.47 85.75 48.55 89.36 75.63 87.89 69.26 83.37 73.71 87.96 87.49 85.31
MultiRole-R1 SelfConsis SFT+GRPO(RS) 95.55 89.58 49.06 91.78 75.84 96.54 71.48 90.55 75.12 92.98 89.79 88.45
@@
Zero-shot CoT 85.01 68.06 36.82 79.18 73.63 72.20 75.05 71.83 75.85 83.09 85.58 70.68
Self-Refine 90.42 80.13 49.04 69.40 76.48 83.22 71.28 78.22 76.55 82.31 84.73 80.24
Role-Play 91.18 81.87 49.90 75.73 77.16 75.30 67.41 70.74 75.71 79.05 91.50 76.60
More think 94.57 80.67 41.60 84.04 79.36 83.33 75.90 76.81 79.36 81.77 88.76 80.94
SelfConsis SFT 94.40 75.06 50.98 81.04 81.45 71.34 76.08 73.65 81.50 77.60 91.61 91.62
SelfConsis SFT+DPO 94.98 67.21 51.33 65.03 81.88 42.32 75.82 71.71 79.82 69.41 90.92 68.69
SelfConsis SFT+GRPO 95.98 86.88 51.73 90.33 83.50 89.42 75.65 84.92 81.19 89.64 91.87 86.36
MultiRole-R1 SelfConsis SFT+GRPO(RS) 97.50 90.17 53.98 91.32 86.00 92.89 76.50 89.08 82.00 91.61 93.43 87.24
@@
Zero-shot CoT 91.71 70.10 42.13 60.99 72.29 76.68 73.40 57.43 80.81 57.84 85.41 70.49
Self-Refine 88.93 53.07 45.25 85.00 70.64 49.60 69.40 48.16 69.22 50.99 84.91 82.31
Role-Play 89.77 41.58 47.57 54.89 70.67 47.83 70.26 50.49 72.98 48.17 93.58 81.93
More think 95.18 74.20 43.39 78.98 78.26 72.45 75.10 68.44 81.20 73.07 90.02 75.07
SelfConsis SFT 94.05 74.02 50.32 77.07 78.39 68.35 75.96 72.78 81.00 73.50 91.62 70.23
SelfConsis SFT+DPO 94.21 74.83 50.68 74.14 78.09 50.31 76.10 69.58 80.45 64.87 91.84 65.32
SelfConsis SFT+GRPO 95.91 85.47 51.37 87.74 79.82 86.84 77.83 80.36 81.19 86.40 91.97 85.98
MultiRole-R1 SelfConsis SFT+GRPO(RS) 96.98 88.15 51.72 89.88 81.95 89.09 77.95 83.84 82.10 87.93 94.98 86.93
@@ 4.1 Datasets
We train our model on 3 subjective tasks: ambiguous question answering (BBQ) by Parrish et al. (2022), opinion-based QA (GlobalOpinionQA) by DURMUS et al. (2024), and ethical dilemma (ETHICS) by Hendrycks et al. (2021). To evaluate the effectiveness and generalizability of our approach, we test on 4 additional datasets: cultural natural language inference (CALI) by Huang & Yang (2023), commonsense reasoning (CSQA) by Talmor et al. (2019), and mathematical reasoning (GSM8K) by Cobbe et al. (2021). We also evaluate our method on the more advanced math reasoning task AIME 2024111https://huggingface.co/datasets/Maxwell-Jia/AIME_2024) and present results in Section 5. Specifically, for the out-of-domain (OOD) data, CALI consists of subjective questions, while CSQA, GSM8K and AIME 2024 consist of objective questions. It is worth noting that among these benchmarks, GLOQA and CALI have role-dependent ground truths, whereas the others rely on a single ground truth for all roles.
@@ 4.2 Baselines
We first incorporate the following in-context learning (Brown et al., 2020) settings: (1) Zero-Shot CoT (Kojima et al., 2023; Wei et al., 2023), (2) Role Playing Prompting (Kong et al., 2024) and (3) Self-Refine Prompting (Madaan et al., 2023).
@@
As observed by Muennighoff et al. (2025), extending the reasoning chain length can further enhance the reasoning capabilities of o1-style models. In MultiRole-R1, this is achieved by suppressing the end-of-thinking token and appending a continuation string (e.g., “wait, I need to think from {role}’s perspective”) to encourage extended reasoning from a different role perspective. In the more think baseline, we employ a reasoning length three times longer than regular think, as it offers a balance between efficiency and accuracy based on our pilot analysis.
@@
We perform supervised finetuning on the base model on the self-consistency filtered dataset of size 2,700. Our SFT training and evaluation are conducted via Llama-Factory (Zheng et al., 2024).
@@
We introduce another SFT+RL pipeline as an comparison to MultiRole-R1. In this setting, DPO is applied to the self-consistent SFT model, using ground-truth-hinted role answers as positive samples and inconsistent role answers as negative samples.
@@ 4.3 Models
Our experiment is performed on DeepSeek-R1 series (DeepSeek-AI et al., 2025) including R1-Distill-Qwen-7B, R1-Distill-Llama-8B and R1-Distill-Qwen-14B models, and Qwen3-8B (Qwen Team, 2025) with reasoning mode.
@@ 4.4 Metrics
Taking into account the subjective nature of role-based reasoning where the ground truth for subjective questions vary across different roles, we adopt two different perspective merging strategies during evaluation. The ground truth of the dataset $\mathcal{G}$ is defined as a role ($r$) ground-truth ($g$) pair, defined as $\mathcal{G}=\{r_{i}:g_{i}\}_{i=1}^{|\mathcal{G}|}$. If not specified, accuracy refers to pass@1 accuracy.
(1) Divergent Merging: for tasks such as CALI and GLOQA, each role $i$’s answer $a_{i}$ is compared with the corresponding ground truth $g_{i}$, where the divergent accuracy is given by: $Acc_{div}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[a_{i}=g_{i}].$
(2) Convergent Merging: datasets like BBQ, ETHICS, CSQA, GSM8K and AIME 2024 have answers invariant with role-perspectives. We aggregate different role’s answer to obtain a consensus, and then compare it to the ground truth: $\hat{a}=argmax\sum_{i}\mathbbm{1}(a_{i}=\hat{a}),Acc_{con}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\hat{a}_{i}=g_{i}].$
@@
To quantify the diversity of model-generated reasoning, we design a composite metric that captures multiple level of linguistic diversity, including lexical, structural and discourse domains. Inspired by prior work on lexical and entropy-based diversity in natural language generation (Li et al., 2016; Tanaka-Ishii & Aihara, 2015), our metric is a weighted sum of eight complementary diversity signals, including lexical, token entropy, sentence length, sentence pattern, adjacent sentence, Yule’s K, distinct N-gram and function word diversity. Formal definition of the diversity metrics can be found in Appendix F. Formally, we express the final combined diversity score as:
$\begin{split}D_{final}=\sum_{i}\omega_{i}D_{i},~D_{i}\in\{D_{lex},D_{ent},D_{pat},D_{bi},D_{len},D_{adj},D_{yule},D_{func}\}.\end{split}$ (4)
The choice of the weighting is illustrated in Section 5.
@@ 4.5 Main Results
MultiRole-R1 is effective and generalizable to unseen objective tasks.
Table 1 shows that MultiRole-R1 outperforms almost all baselines, by an average of 10.6% accuracy gain and 18.3% diversity gain. By training on subjective questions solely, MultiRole-R1 shows a 14.1% accuracy gain in in-domain (ID) tasks, and a 7.64% improvement in out-of-domain (OOD) objective and subjective tasks compared with zero-shot CoT. Notably, our method even yields a 5.78% accuracy gain on the unseen, challenging math reasoning dataset AIME 2024, which will be further illustrated in Section 5. This demonstrates the effectiveness and generalizability of our method.
On-policy RL is more suited for diversity enhancement.
@@
We found that on-policy algorithm like GRPO leads to more accurate (+2.44%) and more diverse responses than off-policy RL like DPO (+19.73%). We attribute this to a fundamental mismatch between DPO’s training format and the nature of our task. Subjective questions often have equally valid answers, while the positive-negative pair format of DPO cannot effectively model the diverse, equally valid ground truths inherent to subjective questions.
Perspective diversity is the primary driver of performance.
In the average of 10.6% accuracy gain, MultiRole-R1, 7.5% is contributed by perspective diversity enhancement in SFT, and 3.1% is contributed by GRPO with token-level diversity reward shaping. This verifies the cruciality of enhancing perspective diversity, which is primarily optimized during SFT.
Accuracy gains come from diversity, not verbosity.
Surprisingly MultiRole-R1 also leads to higher reasoning efficiency. Tables 912 in the appendix show that the average response lengths for SFT, SFT+GRPO, and MultiRole-R1 are 1572.9, 849.5, and 657.8 words. This appears to contradict to recent test-time scaling findings (Muennighoff et al., 2025; Ballon et al., 2025), which suggest that longer reasoning often leads to higher performance. This is further discussed in Section 5.
5 Analysis
Accuracy-Diversity Correlations
@@ 5 Analysis
Table 2: A per-task comparison of the correlation coefficient (r, in %) between accuracy and diversity, versus accuracy and length.
Model BBQ GLOQA ETHICS CALI CSQA GSM8K
Acc-Div Acc-Len Acc-Div Acc-Len Acc-Div Acc-Len Acc-Div Acc-Len Acc-Div Acc-Len Acc-Div Acc-Len
R1-Distill-Qwen-7B 94.2 65.5 45.5 53.6 89.2 89.8 98.6 90.6 92.4 50.5 58.4 23.1
R1-Distill-Llama-8B 89.0 62.1 48.0 56.9 82.4 76.0 83.6 65.2 85.6 63.8 64.9 77.1
R1-Distill-Qwen-14B 82.0 -16.0 59.0 35.8 60.5 60.1 77.2 23.6 57.4 61.3 79.2 67.8
Qwen3-8B 76.0 44.3 65.4 63.4 89.3 86.5 73.0 27.4 82.5 43.5 33.4 57.3
@@
Our analysis reveals a positive correlation between accuracy and diversity, as suggested in Table 2. This relationship is reinforced by a strong per-task correlation between diversity and accuracy (0.736 on average), which markedly exceeds the correlation with response length (0.554 on average). According to Table 9 to 12, SFT responses are often the longest in length, but they are less factually correct. We also observed a tendency to repeat answers in SFT model outputs, which is likely caused by reward hacking during the single-verifiable reward during post-training. These results suggest that performance gains in LRM subjective reasoning are driven by a scaling law of diversity, rather than superficial verbosity. It also demonstrates that MultiRole-R1 is able to improve reasoning efficiency. One possible explanation is that optimizing for diversity can serve as a useful inductive bias, enabling the model to explore a broader solution space and discover more accurate, perspective-aligned answers in subjective tasks.
Figure 3: Qualitative example of the 32 most frequent roles in the training data generated by LRMs.
Qualitative Example of Roles
We leverage a prompt-based method and let LLM generate roles pertinent to the context of the question. In the prompt, we specify that the role perspectives need to be contrastive. We also manually scrutinized and modified the generated roles. In total, there are 968 distinct roles, enhancing the perspective diversity of the LRM. As shown in Figure 3 these roles are generated from the train set and cover broad categories, including different moral philosophies, nationalities, identity groups, and specific individuals pertinent to the questions. Figure 3 presents the most frequent roles, where the circle radius is proportional to the occurring frequency. A more detailed visualization is presented in Figure 5.
Filtering Method
Our methodology compares two data sampling strategies for Supervised Fine-Tuning (SFT): an unsupervised self-consistency approach and a supervised method that leverages ground-truth agreement. Table 3 reports the test performance on equal-sized datasets generated by each strategy. The results indicate that while self-consistency filtering can yield slightly lower accuracy in some cases, its performance is broadly comparable to, and can even outperform supervised filtering. This highlights that using noisy, unsupervised data is sufficient and viable for our tasks (Wang et al., 2025c).
BBQ GLOQA CALI ETHICS CSQA GSM8K
(R1-Distill-Qwen-7B)
Const. Filter 85.55 47.13 67.35 67.45 66.88 80.62
GT Filter 88.40 45.55 65.95 68.44 68.45 80.63
(R1-Distill-Llama-8B)
Const. Filter 89.69 48.17 72.05 70.56 70.86 83.30
GT Filter 87.44 49.29 69.27 72.15 71.06 84.20
(R1-Distill-Qwen-14B)
Const. Filter 94.40 50.98 76.98 81.45 80.50 91.61
GT Filter 94.88 52.29 76.28 81.57 80.79 91.28
(Qwen3-8B)
Const. Filter 94.05 50.32 75.96 78.39 81.00 91.62
GT Filter 94.80 51.07 76.15 80.19 82.13 91.85
Diversity Weighting
Equation (4) shows the composition of diversity score. Because the target application may value different forms of diversity, we adopt equal weight to avoid bias toward any single factor. The final score is the average of all metrics, and we assess its validity by measuring agreement with human annotators. Table 4 shows the alignment scores between diversity score and human ratings. We manually label 60 outputs from different settings, on a diversity scale of 1 to 10. Results show that the overall diversity and the human rating are highly aligned. Table 4 shows that our diversity metric has a high alignment score with human ratings, showcasing the reasonableness of the diversity weighting.
Model BBQ GLOQA CALI (OOD) ETHICS CSQA (OOD) GSM8K (OOD) Alignment
@@
Math Reasoning Generalization
Model R-D-Qwen-7B R-D-Llama-8B R-D-Qwen-14B Qwen3-8B
@@
To assess our framework’s generalizability to complex mathematical reasoning, we evaluated MultiRole-R1 on the AIME 2024 benchmark, which is considerably more challenging than GSM8K. Notably, our method demonstrates a consistent OOD performance gain on AIME 2024, achieving a notable 5.78% average accuracy gain. This suggests the diversity training on subjective questions is transferable to the objective domain, demonstrating that promoting perspective and token-level diversity is a general that could be applied to other domains.
Effect of Diversity Reward Shaping
Figure 4: Comparison of Pass@k accuracy of R1-Distill-Qwen-7B on GLOQA dataset, w/ and w/o diversity reward shaping.
As pass@k accuracy is a direct indicator of reasoning path diversity (Song et al., 2025a), we examine whether the reward shaping broadens the search space of CoT in the GRPO training on the challenging subjective dataset, GLOQA, where k = 5. As shown in Figure 4, training with only a verifiable reward leads to a decreasing pass@k, indicating a convergence towards homogeneous reasoning paths. In contrast, incorporating a diversity reward yields a steady increase in pass@k, confirming that this approach successfully expands the solution search space and is a crucial component of our framework.
6 Related Work
Subjective Tasks and LLM Role-Playing
Subjective tasks lack a single ground truth; answers can shift with perspective or context (Wang et al., 2025b; Jentzsch & Kersting, 2023; Wu et al., 2024a). Examples include culture-related QA (Huang & Yang, 2023; DURMUS et al., 2024; Huang et al., 2025; Liu et al., 2025a), subjective language interpretation (Jones et al., 2025), ethical QA (Hendrycks et al., 2021), and creative QA (Lu et al., 2024). LLM role-playing, including multi-role debate, is a common method: systems simulate assigned personas, from real figures to fictional characters (Shao et al., 2023; Wang et al., 2024a; Du et al., 2023; Liang et al., 2024; Chen et al., 2024; Li et al., 2025), which has been shown to diversify reasoning paths (Naik et al., 2024; Wang et al., 2024c). In this work, we introduce perspective diversity in model training, enabling the model to think about different answers that are equally valid to subjective questions. (Wang et al., 2025b).
Diversity-enhanced Training
Diversity-enhanced training has been recognized as effective in promoting LLM reasoning ability. One category of diversity-enhanced training is finetuning the model on the a set of synthesized diverse reasoning chains (Peng et al., 2025; Zelikman et al., 2022; Chen et al., 2023; Lv et al., 2024). Recently, reinforcement learning with verifiable rewards (RLVR) is widely discussed by the research community due to the diversity collapse issue (Song et al., 2025b; Dang et al., 2025; Yue et al., 2025). Some works use diversity mainly to improve exploration efficiency (Hong et al., 2018; Cheng et al., 2025; Zheng et al., 2025; Dang et al., 2025; Song et al., 2025a). The other works take diversity as a regularization term or objective (Masood & Doshi-Velez, 2019; Yan et al., 2025; Zhou et al., 2022). So far, these methods focus on objective reasoning, and we are the first to apply diversity-enhanced reinforcement learning on subjective questions.